Paragraph Question Three a) The table shows the number of people employed in10 small factories and their annual sales in thousands Firm No.of workers 60 25 65 48 35 28 75 50 45 85 Total sales 250 160 108 151 85 96 151 273 241 290 Calculate the Spearman Rank correlation coefficient and comment. A BCDE F G HK b) In Kenyas industrial sector 10 new products have been developed. The cost of developing these products and the corresponding sales and revenue are shown in the following table Product Development Cos(sh. M) 24 23 25 14 21 22 21 18 192 Sales Revenue (Sh. M) 47 49 57 33 43 4 46 35 40 44 (0 Fit the regression of sales revenue in development cost and estimate sales revenue given a development cost of 22.8M (sh.) () Compute the coefficient of determination and comment. c) The amount of cotton grown (Kg.000) in a country during 2013 to 2016 is shown in the following table: Quarter Year 2013 2014 2015 2016 25 22 24 27 28 23 25 29 30 25 27 31 32 27 29 33 By using moving average of suitable order, compute the trendvalues Prtsc Inser F9 禮
Answer #1

a)

SpearMan’s Rank Correlation

Steps

1) Calculate Ranks (in descending order) for both No Of workers and Total Sales and name them as X and Y respectively. Total of d will be zero

2) Then we calculate difference (d) for each Firm

3) Then we sum of D Square

Rank Firm No Of Workers Total Sales X d ( X-Y) d Squared 4 160 108 151 10 6.5 0.5 0.25 35 10 4 28 -4.5 151 273 241 290 6.5 20.25 4 Total 93.5

Spearman Rank Correlation is given by:

2 ρ=1-n(n2-1)

Hence,

Coefficient = 0.433

The value suggests mild positive correlation

b)

Regression Output

Regression Statistics MultipleR R Square Adjusted R Square Standard Error Observations 0.9213 0.8488 0.8299 2.8365 10 ANOVA MS Significance F Regression Residual Total 361.24361.24 44.90 0.00 64.36 425.60 8.05 9 Coefficients 2.80 1.98 Standard Error 6.18 0.30 P-value Lower 95% Upper 95% 17.06 2.66 tstat Intercept 0.45 0.66 11.45 6.70 0.00 1.30

Regression Equation

Sales Revenue = 2.8 + 1.98 * Development Cost

When Development Cost = 22.8

Sales Revenue = 2.8 + 1.98 * 22.8 = 47.96

ii)

Coefficient of Determination = 0.8488 or 84.88%

c)

Steps:

  1. First we calculate 4 months moving Average.
  2. Then we calculate Centred Average of these 4 months moving Average.

Year QuarterCotton Growth 4QMA Centred Average 2013 01 2013 02 2013 Q3 2013 Q4 2014 01 2014 02 2014 Q3 2014 04 2015 01 2015 Q2 2015 Q3 2015 Q4 2016 01 201602 201603 2016 04 25 24 27 28 23 25 29 30 25 27 31 32 27 29 24.50 25.25 25.50 25.75 26.25 26.75 27.25 27.75 28.25 28.75 29.25 29.75 30.25 24.88 25.38 25.63 26.00 26.50 27.00 27.50 28.00 28.50 29.00 29.50 30.00

QuarterCotton Growth Centred Average 01 02 03 Year 25 2013 24 01 02 03 28 23 25 29 30 24.875 25.375 25.625 26 26.5 2014 01 02 03 27.5 28 28.5 29 29.5 30 2015 31 32 01 02 03 2016 29

Cotton Growth Chart (2013-16) Cotton GrowthCentred Average 35 30 25 20 15 10) 01 2 3 4 01 02 3 Q4 01 2 3 Q4 01 3 04 2014 2015 2016